OpenStudy (anonymous):

integral of 5/(9+4r^2) dr

OpenStudy (luigi0210):

Before suggesting anything, what methods do you know for integration?

OpenStudy (anonymous):

u-sub hmm i am aware of trig and partial fractions but our teacher hasn't taught

OpenStudy (anonymous):

i just get to 5 int (1/(9+r^2)) then i am stuck, i can't find a way to u-sub or "add zero" to make the fraction better

OpenStudy (solomonzelman):

\[\LARGE\color{blue}{ \int\limits_{ }^{ } \frac{5}{4r^2+9}~~dr}\] \[\LARGE\color{blue}{ \frac{5}{9}\int\limits_{ }^{ } \frac{1}{\frac{4r^2}{9}+1}~~dr}\] and then substitute `u` for `2r/3`

OpenStudy (luigi0210):

I'll wait for Solomon to reply, he might have it~

OpenStudy (solomonzelman):

I am not very good at this though, because the heights math I took was trig.

OpenStudy (solomonzelman):

I can finish if you want to .

OpenStudy (solomonzelman):

the highest...

OpenStudy (anonymous):

yes please But what made you decide to divide the denominator by 9?

OpenStudy (solomonzelman):

I am taking the constant out, which is 5/9 in this case.

OpenStudy (solomonzelman):

I don't want to take the responsibility..... @luigi, please -:( :)

OpenStudy (anonymous):

hmm using 2r/3 for you i get 15/18 sintegral of 1/(1+u2)?

OpenStudy (luigi0210):

You're one the right path so far Solomon, now just use \(\Large \int \frac{1}{1+u^2}~du=tan^{-1}u+C\)

OpenStudy (anonymous):

AH! yes haha i forgot about the inverse function!

OpenStudy (solomonzelman):

I was doing it correctly, and I am getting \[\LARGE\color{blue}{ \frac{5}{9}\int\limits_{ }^{ } \frac{1}{\frac{4r^2}{9}+1}~~dr}\]